A proton accelerates from rest in a uniform electric field of 680 N/C. At some later time, its speed is 1.48 x 106 m/s. (a) Find the magnitude of the acceleration of the proton. 6.515E10 m/ls2 (b) How long does it take the proton to reach this speed? 1.84195-5X Use the acceleration you calculated in part (a) of this problem to find the time required for the proton to reach the specified speed. (c) How far has it moved in that interval? 16.8 (d) What is its kinetic energy at the later time? 1.2046E-15X Use the mass and final speed of the proton to calculate its final kinetic energy.Please show work
Answer #1

at 14 d )

EDIT:

(d) KE = 1/2*m*v^2

= 1/2*(1.67*10^-27)(1.48*10^6)^2

= 1.83*10^-15 J